2yzdx-2xzdy-(x^2-y^2)(z-1)dz=0
1
Expert's answer
2020-12-28T13:09:51-0500

"2yzdx-2xzdy-(x^2-y^2)(z-1)dz=0........................(1)"

First we take "z" to be constant. Then the equation reduces to

"2yzdx-2xzdy=0"

Divide through by "2xyz"

"\frac{dy}{y}-\frac{dx}{x}=0"

Integrating, we obtain;

"\ln(\frac{y}{x})=c\\
y=xc_1"

We assume that the solution of Eqn(1) is

"y=x\Psi(z)"

Now, let "x=\alpha" and, therefore,

"y=\alpha\Psi(z) \text{ is a solution of eqn(1).}"

Since "x=\alpha,dx=0" . Put this into eqn(1)

"2\alpha zdy+(\alpha^2-y^2)(z-1)dz=0"

Divide through by "z(\alpha^2-y^2)"

"\frac{2\alpha}{\alpha^2-y^2}dy+\left(1-\frac{1}{z}\right)dz=0"

Integrate through

"\ln|\frac{\alpha+y}{\alpha-y}|-\ln|z|+z=\ln|c| \text{ where } c \text{ is a constant.}"

"\implies \frac{\alpha+y}{\alpha+y}=cze^{-z}"

But "y=\alpha \Psi," therefore,

"\frac{\alpha(1+\Psi)}{\alpha(1-\Psi)}=\frac{1+\frac{y}{x}}{1-\frac{y}{x}}=\frac{x+y}{x-y}=cze^{-z}"

Hence the solution is

"(x+y)e^z=c(x-y)z"